# Radial Basis Function Techniques for Neural Field Models on Surfaces

### v1

@@ 4.3 Convergence on the unit square
We evaluate convergence of RBF-QF and neural field simulations on the unit square, using randomly chosen quadrature nodes and triangulations (See Section 4). The test function is $f(x,y)=T_{5}(2x-1)T_{4}(2y-1)$, a product of Chebyshev polynomials (Figure 4.2 C). Using $\varphi(r)=r^{3}$ and stencil size $k=21$, we compute quadrature weights. As shown in Figure 4.2A,C, errors decrease rapidly with increasing appended polynomial degree, often faster than $\mathcal{O}(h^{\text{deg}})$. Panel B shows similar convergence for the simulation of the neural field solution, which we now describe.
Figure 4.2: A Convergence of RBF-QF quadrature for the function $T_{5}(2x-1)T_{4}(2y-1)+1$ on the unit square. We use the RBF $\varphi(r)=r^{3}$, stencil size $k=21$, and append polynomials terms up to the specified degree (denoted by color). B Convergence of a neural field simulation using the same quadrature rules (see main text for details). C Same quadrature convergence test as in A, but on regular triangular meshes. D Neural field solution at times $t=2n\pi$ ($n\in\mathbb{Z}$), forming a stable bump whose center traces a circular path (shown in white).
@@
which is indeed a solution to eq. 21 on $(t,\bm{x})\in[0,1/10]\times D_{2\pi}$. Here $f$ is defined in eq. 20, and the weight kernel is $w(\bm{x},\bm{y})=\text{Gauss}(\bm{x},\bm{y};\sigma=1/40)$ (See eq. 19). These expressions are used to construct a consistent forcing function. Figure 4.2D shows a snapshot of the resulting traveling bump solution and its circular trajectory, while panel B reports convergence results for different polynomial degrees. In each case, the observed order of accuracy exceeds the expected $\mathcal{O}(h^{\text{deg}})$ rate.
Figure 4.3: A. Convergence of surface quadrature on the torus $D_{T}$ using the test function $f(x,y,z)=\sin(7x)+1$. In all cases, observed rates exceed the degree of appended polynomials. B. Convergence of a neural field simulation using the same surface quadrature rule. C. Torus surface colored by the quadrature weights for $k=12$ and polynomial degree 2. Due to mesh regularity, weights are constant along $\theta$ and larger in regions of positive curvature where triangles are bigger. D Histogram of the quadrature weights shown in panel C.
@@ 4.4 Convergence on a Torus
To test the surface RBF-QF algorithm directly, we use the function $f(x,y,z)=\sin(7x)+1$ on the torus $D_{T}$, with $\varphi(r)=r^{3}$, stencil size $k=21$, and an a polynomial basis appended up to a specified degree. While we lack a direct mesh radius on the surface, the nodes are well spaced, so we use $n^{-1/2}=\mathcal{O}(h)$ as a proxy measure of resolution [49, 51]. Figure 4.3A shows that for all polynomial degrees tested, convergence exceeds the expected rate $\mathcal{O}(n^{-\text{deg}/2})$.
@@
This mirrors the previous solution but is defined on the surface domain $D_{2\pi}$ (in $\varphi$-$\theta$ coordinates), with the weight kernel adjusted by the inverse Jacobian of the torus parametrization ($[R+r\cos\theta]^{-1}$). Although the analytical solution is identical, the quadrature weights differ due to the use of surface-specific integration techniques.
Results are shown in Figure 4.3. Panel A shows quadrature convergence on the torus using the test function $f(x,y,z)=\sin(7x)+1$ with relative error decreasing as $n^{-1/2}$ and higher decay rates for higher-degree appended polynomials. Panel B reports convergence for the manufactured neural field solution, again showing high-order accuracy. Panel C visualizes the quadrature weights across the torus surface for $n=32000$, while D presents a histogram of these weights. Weights are constant along lines of constant $\theta$ due to the symmetry of the node placement; their distribution reflects both curvature effects and histogram binning.
@@ 5.1 Labyrinthine patterns on a deformed sphere
$w(\bm{x},\bm{y})$ $=A_{e}\,\text{Gauss}(\bm{x},\bm{y},\sigma_{e})-A_{i}\,\text{Gauss}(\bm{x},\bm{y},\sigma_{i}),$ (24)
$f(u)$ =$=\begin{cases}0,&u<0.06,\\ p(u),&0.06\leq u<0.54,\\ 1,&u\geq 0.54,\end{cases}$ (25)
where $A_{e}=5$, $A_{i}=5$ $\sigma_{e}=0.05$, $\sigma_{i}=0.1$. Although $A_{e}=A_{i}$, the kernel satisfies $w(\mathbf{x},\mathbf{x})>0$ since the Gaussian is normalized – its peak is higher for smaller $\sigma$ – producing the desired lateral inhibitory property. The function $p(u)$ is a $9$th-order polynomial chosen to ensure that $f$ is four times continuously differentiable, yielding a smooth spline approximation to the Heaviside function $H(u-0.3)$, remaining constant outside a narrow transition interval centered at the threshold.
@@
and is visualized in Figure 5.1 for $\gamma=0,0.4,0.8$.
Figure 5.1: Effect of Surface Deformation on Labyrinthine Patterns. Labyrinthine patterns on a deformed spherical surface $D_{\gamma}$, with increasing deformation parameter from left to right: A. $\gamma=0$ (perfect sphere), B. $\gamma=0.4$ (moderate deformation), C. $\gamma=0.8$ (flattened, blood-cell-like shape). The top row shows the initial unstable, cross-shaped activity pattern. The middle and bottom rows show the resulting spatiotemporal patterns from two viewpoints: with the north and south poles facing the viewer, respectively. Yellow indicates high neural activity ($u>0$), green denotes baseline activity ($u\approx 0$), and dark green/blue indicates suppressed activity ($u<0$). Surface curvature has a profound effect on the form and complexity of the resulting patterns. Click image to view full animation (Movie S1A, Movie S1B, Movie S1C).
@@
which defines a symmetric, cross-shaped pattern localized in the northern hemisphere. To orient this pattern appropriately, we apply a rigid rotation that maps the north pole $[0,0,1]$ to the point $[0.5,0.3,\sqrt{0.5^{2}+0.3^{2}}]$, aligning the initial condition with the desired surface region before the deformation is applied.
We then simulate the neural field on each deformed surface using the kernel basis function $\varphi(r)=r^{3}$, a stencil size of $k=32$, and append third-degree polynomials for accurate quadrature. Time integration is performed using the Adams–Bashforth 5 method with a fixed time step of $\Delta t=10^{-2}$. The resulting dynamics, corresponding to various values of the surface deformation parameter $\gamma=0,0.4,0.8$, are shown in Figure 5.1. The top row depicts the initial condition, rendered from a viewpoint where the north pole is visible. The middle row shows the activity at time $t=200$ from the same viewpoint, while the bottom row presents the state at $t=200$ with the surface inverted through the $x$-$y$ plane to reveal the south pole.
@@ 5.2 Traveling spot steered by surface bumps
[…]
@@
Figure 5.2: Traveling spot steered by bumps in curvature. A–B: Snapshots from a simulation of a traveling spot on a bumpy sphere. Each snapshot shows two views: the left sphere is colored by the activity variable, and the right is shaded by the synaptic efficacy. The trajectory of the spot up to that point is shown as a red curve hovering above the surface. C: Projection of the bumpy surface into spherical coordinates, with longitude on the horizontal axis, latitude on the vertical axis, and surface elevation encoded by heatmap. Traveling spot trajectory (red) sometimes hugs but also deviates from great-circle tangents (white) aligned spot position in panels A and B, showing how bumps deflect motion. Click image to view full animation (Movie S2).
@@
We probe how surface curvature affects trajectories of such traveling spots by considering a surface that we refer to as a bumpy sphere (Figure 5.2), defined by the equation:
$1+\sum\limits_{\bm{x}_{i}}\text{Gauss}^{\prime}(\bm{x},\bm{x}_{i},1/10)/10=x^{2}+y^{2}+z^{2}$
@@
Figure 5.2A and B each show two snapshots from the same simulation, each depicting two views of the bumpy sphere. The left view is colored by the activity variable, while the right shows synaptic efficacy in grayscale. Both are rotated to center the spot in view, and the grayscale surface includes a red curve traving the trajectory of the spot up to that time. Figure 5.2C presents the surface in spherical coordinates, with latitude and longitude along the vertical and horizontal axis, and color representing radial elevation. The red curve shows the full trajectory of the traveling spot. If the surface were perfectly spherical, the spot would follow a straight path along a great circle, akin to what was found on the planar case [13]. Instead, we observe consistent deviations in the trajectory that arise from geometric inhomogeneities. Two white curves show tangent great circles at times shown in panels A and B, illustrating how the trajectory is deflected by geometric inhomogeneities. This reenforces the point that surface irregularities creates a similar potential surface which shape the dynamics of evolving spatiotemporal solutions, akin to those created in neural fields on flat domains with weight inhomogeneities [46, 36].
@@ 5.3 Labyrinthine patterns on a human cortex
Figure 5.3: Labyrinthine dynamics on a cortex. Here we show a simulation on the left hemisphere of a realistic human cortex using a laterally inhibitory weight kernel that favors the formation of labyrinthine patterns. A: Snapshots at evenly spaced early time points show the initial stripe of activity breaking into multiple labyrinth-like corridors that begin to grow and evolve. B–D: Final-time snapshots from different viewing angles highlight distinct cortical regions: B. frontal lobe, C. parietal lobe, and D. limbic cortex/medial surface. Click image to view full animation (Movie S3).
Finally, we build a simulation showcasing the evolution of a neural field on a cortical surface extracted from human data (Figure 5.3). See Code Availability for a link to the repository containing this mesh, representing the left hemisphere of a human cortex. It was generated using MNE-Python, which integrates FreeSurfer’s anatomical reconstruction pipeline to segment T1-weighted MRI scans and extract detailed cortical surfaces. The resulting triangulated mesh captures the geometry of the pial surface, including sulci and gyri, enabling anatomically realistic neural field simulations. We use the same laterally inhibitory weight function defined in eq. 24 with $A_{e}=A_{i}=5$, $\sigma_{e}=3$, $\sigma_{i}=6$, along with the geodesic distance metric and parameters from Section 5.1. The initial condition consists of a band of activity, which rapidly fragments into localized regions of varying size shaped by the surface curvature. The resulting dynamics include both stationary spots and labyrinthine corridors that tend to follow gyri and sulci—regions aligned with locally minimal curvature.
The mesh used here is visually compelling and anatomically detailed, though originally intended for visualization rather than numerical simulation. It exhibits some geometric irregularities – including a small duplicated region, uneven node density, and inconsistencies in surface normals. This leads to rapidly oscillating quadrature weights, including some large negative values. Nonetheless, the resulting dynamics reveal rich qualitative structure that underscores the potential of this framework and motivates further refinement and study.
@@ 6 Conclusion
We have presented a high-order, mesh-flexible solver for neural fields on smooth, closed surfaces, using RBF-based interpolation and quadrature. This method is numerically stable, accurate, and only requires a triangulated surface mesh and approximate vertex normals. Unlike spectral methods, which typically require structured domains, our approach—like finite element methods—can be applied to arbitrary geometries. However, it offers additional flexibility in simulating neural fields on domains with complex curvature without requiring a mesh. The sparse quadrature matrices produced by this scheme still have quadratic computational complexity in the number of nodes due to the pairwise interactions involved. This scaling presents a practical bottleneck, making it especially important to achieve high accuracy with fewer nodes—hence the motivation for high-order methods. This tradeoff is crucial when modeling cortex-scale dynamics or pursuing inverse problems where kernel learning is computationally intensive.
Neural field simulation can thus be performed on deformed/bumpy spheres and realistic human cortical surfaces, where geometry significantly shapes activity patterns. Our results mirror phenomena observed in planar neural fields with inhomogeneous coupling—such as wave pinning and curvature-guided propagation—and show that similar principles extend to curved domains.
We also note that it is common to couple neural field equations with differential equations, or to otherwise include spatial differential operators, commonly diffusion [35, 5]. There are a variety of radial basis function finite difference techniques that approximate operators such as surface advection or surface diffusion that pair nicely with this method [43, 54, 39, 56, 45, 27]. Specifically, the computation of quadrature weights and the computation of finite-difference weights requires solving linear systems with the same matrix, giving opportunity for savings (though the order of accuracy will depend on the order of the differential operator). More importantly, these methods will be high-order and geometrically flexible.
Our results illustrate the type of experiments this method enables and suggest many directions for exploration. In Figure 5.1, the gyrus-like ridge appears to steer the wandering labyrinth along low-curvature paths, confining it to one side until repulsion from other corridors forces a transition. Similar behavior is seen on actual gyri in Figure 5.3. It would be interesting to study how the shape and excitatory–inhibitory balance of the kernel contribute to curvature-guided dynamics, how surface geometry steers traveling spots, and whether a critical angle of incidence maximizes deflection.
The bumpy sphere experiment in Figure 5.2 shows that curvature deflects traveling spot trajectories. Its effects on speed or stability remain unclear, though prior work suggests curvature can pin or disrupt waves [9, 19]. In multi-spot simulations, we observed curvature-modulated crowding and spot death [46, 34]. Future work could reduce spot dynamics to effective equations or study curvature-driven pinning and annihilation [11, 41]. These simulations also invite asymptotic approaches, such as interface methods in the high-gain limit [20], perturbation theory for bump drift [34, 11], or even weakly nonlinear analysis for patterns near bifurcations [61].
@@
Our method leverages only surface geometry and admits arbitrary kernels, providing a promising platform for pairing with experimental data to learn connectivity kernels directly. This opens doors to data-driven modeling, model inversion, and further theoretical studies of how cortical geometry sculpts large-scale dynamics.
@@
Repository addresses containing code developed for implementing quadrature and time-stepping are here: quadrature weight generation: https://github.com/shawsa/neural-fields-rbf; node generation: https://github.com/shawsa/min_energy_points; numerical time-integration: https://github.com/shawsa/odeiter. For curved domains, the first order approximation to the geodesic distance via the Fast Marching Algorithm [53] can be found in the MeshLib library: https://meshlib.io/. Precise centers of bumps on the bumpy sphere are given in https://github.com/shawsa/neural-fields-rbf. The realistic cortical mesh was adapted from the MNE python library: https://mne.tools/stable/index.html.

### v2

@@ 4.3 Convergence on the unit square
We evaluate convergence of RBF-QF and neural field simulations on the unit square, using randomly chosen quadrature nodes and triangulations (See Section 4). The test function is $f(x,y)=T_{5}(2x-1)T_{4}(2y-1)$, a product of Chebyshev polynomials (Figure 4 C). Using $\varphi(r)=r^{3}$ and stencil size $k=21$, we compute quadrature weights. As shown in Figure 4A, errors decrease rapidly with increasing appended polynomial degree, often faster than $\mathcal{O}(h^{\text{deg}})$. Panel B shows similar convergence for the simulation of the neural field solution, which we now describe.
Figure 4: A Convergence of RBF-QF quadrature for the function $T_{5}(2x-1)T_{4}(2y-1)+1$ on the unit square. We use the RBF $\varphi(r)=r^{3}$, stencil size $k=21$, and append polynomials terms up to the specified degree (denoted by color). B Convergence of a neural field simulation using the same quadrature rules (see main text for details).
@@
which is indeed a solution to Eq. 20 on $(t,\bm{x})\in[0,1/10]\times D_{2\pi}$. Here $f$ is defined in Eq. 19, and the weight kernel is $w(\bm{x},\bm{y})=\text{Gauss}(\bm{x},\bm{y};\sigma=1/40)$ (See Eq. 18). These expressions are used to construct a consistent forcing function. Figure 4 B reports convergence results for different polynomial degrees. In each case, the observed order of accuracy exceeds the expected $\mathcal{O}(h^{\text{deg}})$ rate. For random meshes, the noise in the error curves is expected: each random placement of nodes slightly alters local point density and interpolation accuracy, and these local variations accumulate in the global error measure, producing the observed fluctuations.
Figure 5: A. Convergence of surface quadrature on the torus $D_{T}$ using the test function $f(x,y,z)=\sin(7x)+1$. In all cases, observed rates exceed the degree of appended polynomials. B. Convergence of a neural field simulation using the same surface quadrature rule. C. Torus surface colored by the quadrature weights for $k=12$ and polynomial degree 2. Due to mesh regularity, weights are constant along $\theta$ and larger in regions of positive curvature where triangles are bigger. D Histogram of the quadrature weights shown in panel C.
@@ 4.4 Convergence on a Torus
To test the surface RBF-QF algorithm directly, we use the function $f(x,y,z)=\sin(7x)+1$ on the torus $D_{T}$, with $\varphi(r)=r^{3}$, stencil size $k=21$, and an a polynomial basis appended up to a specified degree. While we lack a direct mesh radius on the surface, the nodes are well spaced, so we use $n^{-1/2}=\mathcal{O}(h)$ as a proxy measure of resolution [52, 54]. Figure 5A shows that for all polynomial degrees tested, convergence exceeds the expected rate $\mathcal{O}(n^{-\text{deg}/2})$.
@@
This mirrors the previous solution but is defined on the surface domain $D_{2\pi}$ (in $\varphi$-$\theta$ coordinates), with the weight kernel adjusted by the inverse Jacobian of the torus parametrization ($R+r\cos\theta]^{-1}$). Although the analytical solution is identical, the quadrature weights differ due to the use of surface-specific integration techniques.
Results are shown in Figure 5. Panel A shows quadrature convergence on the torus using the test function $f(x,y,z)=\sin(7x)+1$ with relative error decreasing as $n^{-1/2}$ and higher decay rates for higher-degree appended polynomials. Panel B reports convergence for the manufactured neural field solution, again showing high-order accuracy. Panel C visualizes the quadrature weights across the torus surface for $n=32000$, while D presents a histogram of these weights. Weights are constant along lines of constant $\theta$ due to the symmetry of the node placement; their distribution reflects both curvature effects and histogram binning.
@@ 5.1 Labyrinthine patterns on a deformed sphere
(23) $w(\bm{x},\bm{y})$ $=A_{e}\,\text{Gauss}(\bm{x},\bm{y},\sigma_{e})-A_{i}\,\text{Gauss}(\bm{x},\bm{y},\sigma_{i}),$
(24) $f(u)$ $=\begin{cases}0,&u<0.06,\\ p(u),&0.06\leq u<0.54,\\ 1,&u\geq 0.54,\end{cases}$
where $A_{e}=5$, $A_{i}=5$ $\sigma_{e}=0.05$, $\sigma_{i}=0.1$. Although $A_{e}=A_{i}$, the kernel satisfies $w(\mathbf{x},\mathbf{x})>0$ since the Gaussian is normalized – its peak is higher for smaller $\sigma$ – producing the desired lateral inhibitory property. The function $p(u)$ is a $9$th-order polynomial chosen to ensure that $f$ is four times continuously differentiable, yielding a smooth spline approximation to the Heaviside function H(u0.3)H(u-0.3), remaining constant outside a narrow transition interval centered at the threshold.
@@
and is visualized in Figure 6 for $\gamma=0,0.4,0.8$.
Figure 6: Effect of Surface Deformation on Labyrinthine Patterns. Labyrinthine patterns on a deformed spherical surface $D_{\gamma}$, with increasing deformation parameter from left to right: A. $\gamma=0$ (perfect sphere), B. $\gamma=0.4$ (moderate deformation), C. $\gamma=0.8$ (flattened, blood-cell-like shape). An initial unstable, cross-shaped activity pattern (top) evolves to meandering spatiotemporal patterns (middle: north pole viewpoint; bottom: south pole viewpoint). Yellow indicates high neural activity ($u>0$), green denotes baseline activity ($u\approx 0$), and dark green/blue indicates suppressed activity ($u<0$). Surface curvature has a profound effect on the form and complexity of the resulting patterns. Click image to view full animation (Movie S1A, Movie S1B, Movie S1C).
@@
which defines a symmetric, cross-shaped pattern localized in the northern hemisphere. To orient this pattern appropriately, we apply a rigid rotation that maps the north pole $[0,0,1]$ to the point $0.5,0.3,\sqrt{0.5^{2}+0.3^{2}}]$, aligning the initial condition with the desired surface region before the deformation is applied.
We then simulate the neural field on each deformed surface using the kernel basis function $\varphi(r)=r^{3}$, a stencil size of $k=32$, and append third-degree polynomials for accurate quadrature. Time integration is performed using the Adams–Bashforth 5 method with a fixed time step of $\Delta t=10^{-2}$. The resulting dynamics, corresponding to various values of the surface deformation parameter $\gamma=0,0.4,0.8$, are shown in Figure 6. The top row depicts the initial condition, rendered from a viewpoint where the north pole is visible. The middle row shows the activity at time $t=200$ from the same viewpoint, while the bottom row presents the state at $t=200$ with the surface inverted through the $x$-$y$ plane to reveal the south pole.
@@ 5.2 Traveling spot steered by surface bumps
(25) $\partial_{t}u$ $=-u+\int_{D}w({\mkern 2.0mu\cdot\mkern 2.0mu},\bm{y})q({\mkern 2.0mu\cdot\mkern 2.0mu},\bm{y})f[u({\mkern 2.0mu\cdot\mkern 2.0mu},\bm{y})]\ d\bm{y},\qquad\tau\partial_{t}q=1-q-\beta qf[u].$
@@
Figure 7: Traveling spot steered by bumps in curvature. A–B: Snapshots from a simulation of a traveling spot on a bumpy sphere. Each shows the activity variable (left) and synaptic efficacy (right). The trajectory of the spot up to that point is shown as a red curve hovering above the surface. C: Projection of the bumpy surface into spherical coordinates, with longitude on the horizontal axis, latitude on the vertical axis, and surface elevation encoded by heatmap. Traveling spot trajectory (red) sometimes hugs but also deviates from great-circle tangents (white) aligned spot position in panels A and B, showing how bumps deflect motion. Click image to view full animation (Movie S2).
@@
We probe how surface curvature affects trajectories of such traveling spots by considering a surface that we refer to as a bumpy sphere (Figure 7), defined by the equation:
$1+\sum\limits_{\bm{x}_{i}}\left(\frac{2\pi}{10^{2}}\text{Gauss}(\bm{x},\bm{x}_{i},1/10)\right)\tfrac{1}{10}=x^{2}+y^{2}+z^{2}$
@@
Figure 7A and B each show two snapshots from the same simulation, each depicting two views of the bumpy sphere. The left view is colored by the activity variable, while the right shows synaptic efficacy in grayscale. Both are rotated to center the spot in view, and the grayscale surface includes a red curve traving the trajectory of the spot up to that time. Figure 7C presents the surface in spherical coordinates, with latitude and longitude along the vertical and horizontal axis, and color representing radial elevation. The red curve shows the full trajectory of the traveling spot. If the surface were perfectly spherical, the spot would follow a straight path along a great circle, akin to what was found on the planar case [16]. Instead, we observe consistent deviations in the trajectory that arise from geometric inhomogeneities. Two white curves show tangent great circles at times shown in panels A and B, illustrating how the trajectory is deflected by geometric inhomogeneities. This reenforces the point that surface irregularities creates a similar potential surface which shape the dynamics of evolving spatiotemporal solutions, akin to those created in neural fields on flat domains with weight inhomogeneities [49, 39].
@@ 5.3 Labyrinthine patterns on a human cortex
Figure 8: Labyrinthine dynamics on a cortex. Here we show a simulation on the left hemisphere of a realistic human cortex using a laterally inhibitory weight kernel that favors the formation of labyrinthine patterns. A: Snapshots at evenly spaced early time points show the initial stripe of activity breaking into multiple labyrinth-like corridors that begin to grow and evolve. B–D: Final-time snapshots from different viewing angles highlight distinct cortical regions: B. frontal lobe, C. parietal lobe, and D. limbic cortex/medial surface. Click image to view full animation (Movie S3).
Finally, we build a simulation showcasing the evolution of a neural field on a cortical surface extracted from human data (Figure 8). See Code Availability for a link to the repository containing this mesh, representing the left hemisphere of a human cortex. It was generated using MNE-Python, which integrates FreeSurfer’s anatomical reconstruction pipeline to segment T1-weighted MRI scans and extract detailed cortical surfaces. The resulting triangulated mesh captures the geometry of the pial surface, including sulci and gyri, enabling anatomically realistic neural field simulations. We use the same laterally inhibitory weight function defined in Eq. 23 with $A_{e}=A_{i}=5$, $\sigma_{e}=3$, $\sigma_{i}=6$, along with the geodesic distance metric and parameters from Section 5.1. The initial condition consists of a band of activity, which rapidly fragments into localized regions of varying size shaped by the surface curvature. The resulting dynamics include both stationary spots and labyrinthine corridors that tend to follow gyri and sulci—regions aligned with locally minimal curvature.
The mesh used here is visually compelling and anatomically detailed, though originally intended for visualization rather than numerical simulation. It exhibits some geometric irregularities – including a small duplicated region, uneven node density, and inconsistencies in surface normals. This leads to rapidly oscillating quadrature weights, including some large negative values. However, this is not indicative of an inherent instability of the RBF quadrature method. On other domains, including the flat-domain tests in Section 4.2 and the torus in Section 4.4, the method produces stable, accurate results with only small negative weights, even over long simulations. These observations demonstrate that the method is robust when applied to high-quality, well-distributed node sets. The oscillatory weights in the cortex case stem from the severe irregularity of the publicly available mesh we employed, rather than from the quadrature scheme itself. Nonetheless, the resulting dynamics reveal rich qualitative structure that underscores the potential of this framework and motivates further refinement and study.
@@ 6 Conclusion
We have presented a high-order, mesh-flexible solver for neural fields on smooth, closed surfaces using RBF-based interpolation and quadrature. The method is numerically stable, accurate, and requires only a triangulated mesh and approximate vertex normals. Unlike spectral methods, which rely on structured domains, our approach—like finite element methods—supports arbitrary geometries without requiring element construction. Although the resulting quadrature matrices are sparse, pairwise interactions lead to $\mathcal{O}(n^{2})$ complexity in the number of nodes. This scaling motivates the use of high-order schemes that achieve accuracy with fewer degrees of freedom, especially in cortex-scale modeling or inverse problems involving kernel learning.
Simulations on bumpy spheres and realistic cortical surfaces show how geometry can steer and constrain activity, extending phenomena observed in planar neural fields with inhomogeneous coupling (e.g., wave slowing, deflection, or pinning) to non-Euclidean domains. In Figure 6, a gyrus-like ridge steers a labyrinthine wave pattern along low-curvature paths until repulsion from adjacent corridors forces a transition; similar behavior is observed on real cortical gyri in Figure 8. The bumpy sphere simulation in Figure 7 further demonstrates that curvature can deflect traveling spot trajectories. While its effects on wave speed and stability remain unclear, prior work suggests that curvature can pin or disrupt waves [12]. Multi-spot simulations reveal curvature-modulated crowding and spot annihilation [37], raising broader questions about how surface geometry and the excitatory–inhibitory balance of the kernel interact to steer dynamics, and whether a critical angle of incidence maximizes deflection. These findings motivate future reductions to effective equations and further analysis of curvature-driven pinning, transitions, and stability [14, 45].
Surface differential operators (e.g., diffusion or advection) commonly arise in neural field models with local dynamics [38, 7]. RBF-based finite difference methods provide high-order, geometry-flexible approximations of such operators [47, 56, 42, 58, 48, 30], with the same system matrix used for quadrature weights also yielding finite difference weights—offering computational savings when coupling local and nonlocal dynamics. Other meshfree approaches, such as partition of unity methods [6, 66] and moving least squares [43, 29], share similar flexibility for irregular node layouts and could be adapted to neural field models on surfaces. We focus on RBF-based methods for their direct unification of interpolation, quadrature, and differentiation, noting that recent RBF-FD advances in stabilization and adaptive refinement [9, 26] further expand the toolkit for high-order, geometry-flexible PDE solvers.
@@
Our method depends only on surface geometry and supports arbitrary kernels, offering a promising platform for pairing with experimental data to infer connectivity. This opens the door to data-driven modeling, model inversion, and further theoretical exploration of how cortical geometry shapes large-scale neural activity.
@@
The code used to generate the numerical simulations and figures can be found in the repository www.github.com/shawsa/neural-fields-rbf. For curved domains, the first order approximation to the geodesic distance via the Fast Marching Algorithm [55] can be found in the MeshLib library: meshlib.io. The realistic cortical mesh was adapted from the MNE python library: mne.tools.
Supplementary Material
We present convergence results on the unit square, where collocation nodes are placed on a regular triangular grid in the interior and equally spaced along the boundary. Figure 9A,B shows the node placement and corresponding quadrature weights. Convergence tests are reported in Figure 9C,D for two representative functions: a degree-4 polynomial and a Gaussian. For the polynomial, the quadrature is exact when augmenting the basis with degree-4 polynomials, as expected. The Gaussian is given by
$f(x,y)=\exp\bigg(-10\big[(x-\tfrac{1}{2})^{2}+(y-\tfrac{1}{2})^{2}\big]\bigg)$
decays rapidly to zero near the boundary. Because this function is smooth and effectively supported away from the edges, the quadrature achieves spectral convergence. This behavior is directly analogous to the classical result that the trapezoidal rule attains spectral accuracy for smooth periodic functions [63].
Figure 9: Quadrature on the unit square using a regular triangular grid of collocation nodes, the RBF $\varphi(r)=r^{3}$, and stencil size $k=30$. A. Quadrature nodes colored by weight (negative weights are shown as larger black dots). B. Histogram of quadrature weights. C. Convergence of the quadrature rule for a degree-4 polynomial $x^{3}-y^{4}$. D. Convergence for a Gaussian $f(x,y)=\exp(-10[(x-\tfrac{1}{2})^{2}+(y-\tfrac{1}{2})^{2}])$. Relative error is plotted against $N^{-1/2}$, corresponding to the node length scale.
7 Convergence on other geometries
We next present convergence results for quadrature on curved surfaces, specifically the unit sphere and a Dupin cyclide. For the sphere, collocation nodes are chosen from icosahedral-based point sets [61, 32], which provide nearly uniform coverage. For the cyclide, we consider the ring case with parameters
$a=1,\quad b=0.98,\quad c=0.1983,\quad d=0.5,$
and construct a non-random triangular mesh in the $(\varphi,\theta)$ parameter space.
Figures 10A and 11A illustrate representative meshes. Panels B–D in each figure report relative quadrature error versus $N^{-1/2}$ for three test functions: a constant $f(x,y,z)=1$, a polynomial $f(x,y,z)=x^{3}y^{2}z^{4}+5$, and a trigonometric function $f(x,y,z)=\sin(x)\cos(2y)\cos(3z)$. As predicted, we observe convergence at least has high as the degree of appended polynomial for all three test functions.
Figure 10: Convergence of quadrature on the unit sphere using icosahedral nodes. A. Example of a triangular mesh from icosahedral nodes placement. B-D. Relative quadrature error vs $N^{-1/2}$ (node length scale) for three test functions: a constant (1), a polynomial $x^{3}y^{2}z^{4}+5$, and a trigonometric function $\sin(x)\cos(2y)\cos(3z)$.
Figure 11: Convergence of quadrature on a ring cyclide with parameters $a=1$, $b=0.98$, $c=0.1983$, $d=0.5$, using a non-random triangular mesh in $(\varphi,\theta)$ parameter space. A. Example of the cyclide mesh. B-D. Relative quadrature error vs $N^{-1/2}$ (node length scale) for three test functions: a constant (1), a polynomial $x^{3}y^{2}z^{4}+5$, and a trigonometric function $\sin(x)\cos(2y)\cos(3z)$.